36-303: Sampling, Surveys and Society Summary of Exam 1

Performance on Exam 1 was very good. A histogram appears in Figure 1, and a summary of the score results follow:

Min. 1st Qu.  Median    Mean 3rd Qu.    Max.
54.00   77.00   82.50   82.53   88.25   98.00

Here are approximate letter grades, using a 90/80/70/60 scale:

Score Letter Count
90–100 A  8
80–89 B  12
70–79 C  13
Other  3
Total  39

Figure 1: Histogram of test scores from Exam 1